OpenStudy (anonymous):

Evaluate the following sum:

OpenStudy (anonymous):

\[\sum_{k=0}^{n} k(n-k)\left(\begin{matrix}n \\ k\end{matrix}\right)\]

OpenStudy (zarkon):

use the definition of \({n\choose k}\) to simplify

OpenStudy (zarkon):

and the fact that \[\sum_{k=0}^{m}{m\choose k}=2^m\]

OpenStudy (anonymous):

@zarkon, does it simplify to \[\sum \frac{n!}{(k-1)!(n-k-1)!}\] or am i way off?

OpenStudy (zarkon):

yes

OpenStudy (zarkon):

also notice that \[\sum_{k=0}^{n} k(n-k)\left(\begin{matrix}n \\ k\end{matrix}\right)\] \[=\sum_{k=1}^{n-1} k(n-k)\left(\begin{matrix}n \\ k\end{matrix}\right)\]

OpenStudy (anonymous):

\(n\) is fixed right? so is this the same as \[n\sum\dbinom{n-1}{k-1}\]?

OpenStudy (zarkon):

n is fixed...but you are a little off

OpenStudy (zarkon):

\[n-1-(k-1)=n-k\ne n-k-1\]

OpenStudy (anonymous):

oh yes a am off

OpenStudy (zarkon):

just factor out n-1

OpenStudy (zarkon):

in addition to the n you already factored out

OpenStudy (anonymous):

\[n(n-1)\sum_{}^{} \left(\begin{matrix}n-2 \\ k-1\end{matrix}\right) \ \ \ \ ?\]

OpenStudy (zarkon):

yes

OpenStudy (zarkon):

let \(w=k-1\) then you get \[n(n-1)\sum_{w=0}^{n-2}{n-2\choose w}\]

OpenStudy (anonymous):

the ans is \[n(n-1)2^{n-2} \]:) Thank you!

OpenStudy (zarkon):

yep